Find the volume of the cube bounded by the coordinate planes and the planes X=2 y=2 and z=2 in the first octant
1
Expert's answer
2021-07-14T12:09:30-0400
"V=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{2}dzdydx"

"=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{2}[z]\begin{matrix}
 2\\
 0
\end{matrix}dydx"

"=2\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{2}dydx"

"=2\displaystyle\int_{0}^{2}[y]\begin{matrix}
 2 \\
 0
\end{matrix}dx"

"=4[x]\begin{matrix}
 2 \\
 0
\end{matrix}"

"=8 (units^3)"

"V=8\ cubic\ units"